a) Calculate the pH of a 0.8M solution of HCOOH (Ka = 1.8 x 10-4) in 0.4M Li+ HCOO-.

b) Calculate the pH of a solution of 0.4M NH4Cl and 0.5M NH3 (pKa NH4+ = 9.2).

c) Calculate the pH of a titration of 100 mL 1.5M HCl when 75 mL 1.25M NaOH has been added.

d) What volume of 1.25M NaOH must be added to 100 mL of 1.5M HCl to reach equivalence?
Answer #1

a) HCOOH + 2H Cooli mixture forms a buffer. (Audic buffer) pH of buffer solution i given by pH = pka + log (Salt) plat og ferGiven that (NH3) = 0.5M PnHuel] = 0.4M pke of NHY = 9.2 pka = -log Ka M -962 log Kai -9.2 6.31810 10 = 6.31x10-10 1 .Kor 10 T

lince Phot pot = 14 M M B = 14,004 =14-yit -9.31 c) when v ml of My acid is mixed with Vg ml of Me dbane, the concentration o

During a titration, the eiluhen formule is MV = M V2 Ho, te n ne 10 Na oH THU - Nad to Hy : NaOH M: Molarity rivoluze M = 1.5